OpenStudy (anonymous):

x/2-1=2-x show me the steps

OpenStudy (anonymous):

x = 2

OpenStudy (anonymous):

(x-2)/2=2-x x-2=4-2x 3x=6 x=2

OpenStudy (anonymous):

can you show me more detail?

OpenStudy (anonymous):

(x/2)-1=2-x multiply the whole equation with 2 will get, x-2=4-2x rearranging, 3x=6 x=2

OpenStudy (anonymous):

or you can simplify LHS(left hand side of equation) by taking LCM and simplify like what I did first time.